OpenStudy (anonymous):

A quadratic equation is shown below: x2 + 5x + 4 = 0 Part A: Describe the solution(s) to the equation by just determining discriminant. Show your work. (3 points) Part B: Solve 4x2 -12x + 5 = 0 using an appropriate method. Show the steps of your work and explain why you chose the method used. (4 points) Part C: Solve 2x2 -10x + 3 = 0 by using a method different from the one you used in Part B. Show the steps of your work. (3 points)

OpenStudy (unklerhaukus):

A Quadratic Equation\[ax^2+bx+c=0\] The Quadratic Formula\[x_{1,2}=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\] Factored for of the Quadratic Equation\[(x-x_1)(x-x_2)=0\] start by identifying the coefficients a,b,c Compare \[x^2 + 5x + 4 =0\] With \[ax^2+bx+c=0\] \(a=?\) \(b=?\) \(c=?\) ______________________________ The radicand of the quadratic formula is called the Discriminant \(\Delta =b^2-4ac\) if \(\Delta=0\) there is only one solution to the quadratic equation if \(\Delta>0\) there are two solutions if \(\Delta <0\) the solutions are complex another way of putting this if \(\Delta=0\) there is two equal solutions if \(\Delta>0\) there are two different solutions if \(\Delta<0\) there are no real solutions if \[\Delta \neq 0= d^2 \] the solutions are rational then taking the square root in the quadratic formula will not leave any irrational component \[x_{1,2}=\frac{-b\pm \sqrt\Delta}{2a}=\frac{-b\pm \sqrt d^2}{2a}=\frac{-b\pm d}{2a}\]